# High-temperature series expansion of the dynamic Matsubara spin correlator

### v2

@@ Abstract
The high-temperature series expansion for quantum spin models is a well-established tool to compute thermodynamic quantities and equal-time spin correlations, in particular for frustrated interactions. We extend the scope of this expansion to the dynamic Matsubara spin-spin correlator and develop an algorithm that yields exact expansion coefficients in the form of rational numbers. We focus on Heisenberg models with a single coupling constant $J$ and spin lengths $S\in\{1/2,1\}$. The expansion coefficients up to 12th order in $J/T$ are precomputed on all possible $\sim 10^{6}$ graphs embeddable in arbitrary lattices and are provided under https://github.com/bsbierski/Dyn-HTE. This enables calculation of static momentum-resolved susceptibilities for arbitrary site-pairs or wavevectors. We test our results for the $S=1/2$ Heisenberg chain and on the triangular lattice model. An important application that we discuss in a companion letter is the calculation of real-frequency dynamic structure factors. This is achieved by identifying the high-frequency expansion coefficients of the Matsubara correlator with frequency moments of the spectral function.
@@ I Introduction
Finally, as motivated above, a main application of the Matsubara correlator is its analytical continuation to the real-frequency dynamical spin structure factor [12]. A large set of methods like QMC or pseudo-fermion based diagrammatic approaches [29, 30, 31, 32, 33] produce approximate numerical correlator data on a limited set of points on the Matsubara axis. In such a situation, despite recent advances [34, 35, 36, 37, 38], this analytical continuation is an ill-conditioned and error-prone procedure. For Dyn-HTE, in contrast, the obtained exact expansion can be regrouped in form of a high-frequency expansion (in inverse Matsubara frequency). In our companion work [39] we show that these high-frequency expansion coefficients can be identified with the frequency moments of the (real-frequency) spectral function from which the dynamical structure factor can be reconstructed by standard methods. This means that Dyn-HTE allows to bypass the standard ill-defined analytical continuation procedure.
@@ III Matsubara spin correlator and Dyn-HTE
Our open-source numerical implementation [21] provides the exact polynomial coefficients in Eq. (7) as rational numbers for arbitrary lattices $\mathcal{L}$, all site-pairs $ii^{\prime}$ therein. Currently spin lengths $S\in\{1/2,1\}$ and maximum expansion order $n_{max}=12$ are available.
@@ IV.2 Graph-based evaluation of bond-sums
$c_{ii^{\prime}}^{(n)}(i\nu_{m})$ $=$ $\sum_{g^{(n)}}e(\mathcal{L},i,i^{\prime},g^{(n)})\cdot c_{g^{(n)}}(i\nu_{m}),$ (15)
@@
Finally we need to consider the quantity which we wanted to evaluate on the lattice snippets defined by a selection of edges $\{e_{1},...,e_{n}\}$: The HTE of the Matsubara correlator. We call this HTE the graph evaluation $c_{g^{(n)}}(i\nu_{m})$. In Eq. (15) the correlator expansion on the full lattice is obtained by a sum over all graphs evaluations weighted with the embedding factors. In Eq. (16), the individual graph evaluations are given by a sum $\sum_{\{e_{1},...,e_{n}\}\rightarrow\{\tilde{e}_{1},...,\tilde{e}_{n}\}}$ which associates the multi-set of the graph’s edges to the time-ordered bond operators in $F^{(n)}_{jj^{\prime}}\left(\tilde{e}_{1},...,\tilde{e}_{n};i\nu_{m}\right)$ in all $n!/d[g^{(n)}]$ different ways. Here $d[g^{(n)}]$ yields the graph degeneracy which is the product of factorials of all edge weights.
@@ A.2 Consistency with conventional HTE for uniform susceptibility
As mentioned above, the conventional HTE for equal-time correlators and thermodynamic quantities like the uniform susceptibility $\chi\equiv\sum_{i^{\prime}}\left\langle S_{i}^{z}S_{i^{\prime}}^{z}\right\rangle$ is well established. Here we test Dyn-HTE by reproducing the expansion coefficients of $\chi$ calculated to high orders for many lattices, see below. First, we recover the equal-time correlators from the Matsubara correlators by a frequency sum,
@@
We insert the Dyn-HTE expansion from Eq. (10) on the right-hand side and perform the frequency sums:
@@
As a non-trivial consistency check, we calculate the uniform susceptibility $\chi$ for the triangular lattice Heisenberg $S=1/2$ AFM and reproduce the expansion coefficients provided in Ref. 68 for all orders $n\leq n_{max}=12$. Note that this uses all graph evaluations for graphs embeddable in the triangular lattice with their complete frequency dependence [only a static contribution must survive after the sum in Eq. (44)] and thus constitutes a rather non-trivial check.
For the Kagome lattice, we also reproduce the HTE coefficients for the susceptibility [59]. The $n$-th order coefficient of $\chi$ in (x)(-x) are given by $a_{n+1}(-1)^{n}/[(n+1)!\>4^{n+2}]$ where $a_{n}$ is given in the right column of Tab. I in Ref. 59. There, a misprint must be corrected, $a_{7}=2711296$. This has been also noticed in Ref. 4. For $S=1$ we performed the same check for $\chi$ against the conventional HTE [4].

### v3

@@ Abstract
The high-temperature series expansion for quantum spin models is a well-established tool to compute thermodynamic quantities and equal-time spin correlations, in particular for frustrated interactions. We extend the scope of this expansion to the dynamic Matsubara spin-spin correlator and develop an algorithm that yields exact expansion coefficients in the form of rational numbers. We focus on Heisenberg models with a single coupling constant $J$ and spin lengths $S\in\{1/2,1\}$. The expansion coefficients up to 12th order in $J/T$ are precomputed on all possible $\sim 10^{6}$ graphs embeddable in arbitrary lattices and are provided in a repository. This enables calculation of static momentum-resolved susceptibilities for arbitrary site-pairs or wavevectors. We test our results for the antiferromagnetic $S=1/2$ chain and triangular lattice model. An important application that we discuss in a companion letter is the calculation of real-frequency dynamic structure factors. This is achieved by identifying the high-frequency expansion coefficients of the Matsubara correlator with frequency moments of the spectral function.
@@ I Introduction
Finally, as motivated above, a main application of the Matsubara correlator is its analytical continuation to the real-frequency dynamical spin structure factor [12]. A diverse set of methods like QMC or pseudo-fermion based diagrammatic approaches [29, 30, 31, 32, 33] produce approximate numerical correlator data on a limited set of points on the Matsubara axis. In such a situation, despite recent advances [34, 35, 36, 37, 38], this analytical continuation is an ill-conditioned and error-prone procedure. For Dyn-HTE, in contrast, the obtained exact expansion can be regrouped in form of a high-frequency expansion (in inverse Matsubara frequency). In our companion work [39] we show that the associated expansion coefficients can be identified with the frequency moments of the (real-frequency) spectral function from which the dynamical structure factor can be reconstructed by standard methods. This means that Dyn-HTE allows to bypass the standard ill-defined analytical continuation procedure.
@@ III Matsubara spin correlator and Dyn-HTE
Our open-source numerical implementation [21] provides the exact polynomial coefficients in Eq. (7) as rational numbers for arbitrary lattices $\mathcal{L}$ and all site-pairs $ii^{\prime}$ therein. Currently spin lengths $S\in\{1/2,1\}$ and maximum expansion order $n_{max}=12$ are available.
@@ IV.2 Graph-based evaluation of bond-sums
$c_{ii^{\prime}}^{(n)}(i\nu_{m})$ $=$ $\sum_{g^{(n)}}e(\mathcal{L},i,i^{\prime},g^{(n)})c_{g^{(n)}}(i\nu_{m}),$ (15)
@@
Finally we need to consider the quantity which we wanted to evaluate on the lattice snippets defined by a selection of edges $g^{(n)}=\{e_{1},...,e_{n}\}$: The HTE of the Matsubara correlator. We call this HTE the graph evaluation $c_{g^{(n)}}(i\nu_{m})$. In Eq. (15) the correlator expansion on the full lattice is obtained by a sum over all graph evaluations weighted with the embedding factors. In Eq. (16), the individual graph evaluations are given by a sum $\sum_{\{e_{1},...,e_{n}\}\rightarrow\{\tilde{e}_{1},...,\tilde{e}_{n}\}}$ which associates the multi-set of the graph’s edges to the time-ordered bond operators in $F^{(n)}_{jj^{\prime}}\left(\tilde{e}_{1},...,\tilde{e}_{n};i\nu_{m}\right)$ in all $n!/d[g^{(n)}]$ different ways. Here $d[g^{(n)}]$ yields the graph degeneracy which is the product of factorials of all edge weights.
@@ A.2 Consistency with conventional HTE for uniform susceptibility
As mentioned above, the conventional HTE for equal-time correlators and thermodynamic quantities like the uniform susceptibility $\chi\equiv\sum_{i^{\prime}}\left\langle S_{i}^{z}S_{i^{\prime}}^{z}\right\rangle$ is well established. Here we test Dyn-HTE by reproducing the expansion coefficients of $\chi$ calculated to high orders for many lattices. First, we recover the equal-time correlators from the Matsubara correlators by a frequency sum,
@@
We insert the Dyn-HTE expansion from Eq. (10) on the right-hand side and perform the frequency sum. The numerical values for the sum over even powers of $\Delta_{2\pi m}$ are summarized in Tab. 3 for convenience.
As a first non-trivial consistency check, we calculate the uniform susceptibility $\chi$ for the triangular lattice Heisenberg $S=1/2$ AFM and reproduce the expansion coefficients provided in Ref. 68 for all orders $n\leq n_{max}=12$. Note that this uses all graph evaluations for graphs embeddable in the triangular lattice with their complete frequency dependence [only a static contribution must survive after the sum in Eq. (44)] and thus constitutes a rather non-trivial check.
For the Kagome lattice, we also reproduce the HTE coefficients for the susceptibility [59]. The $n$-th order coefficient of $\chi$ in (x)(-x) is given by $a_{n+1}(-1)^{n}/[(n+1)!\>4^{n+2}]$ where $a_{n}$ is given in the right column of Tab. I in Ref. 59. There, a misprint must be corrected, $a_{7}=2711296$. This has been also noticed in Ref. 4. For $S=1$ we performed the same check for $\chi$ against the conventional HTE [4].
@@
Table 3: Frequency sums needed in Eq. (44).